# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v5

@@ 3 Conclusions
In this work, we have introduced PoseBench, a unified, broadly applicable benchmark and toolkit for studying the performance of DL methods for protein-ligand docking and structure prediction. Benchmarking results with PoseBench suggest that DL co-folding methods generally outperform conventional and DL docking baselines yet remain challenged to predict protein targets containing novel sequences. Further, we find that several DL methods face difficulties balancing the structural accuracy of their predicted poses with the chemical specificity of their induced protein-ligand interactions, highlighting that future methods may benefit from the introduction of physico-chemical loss functions or sampling techniques to bridge this performance gap. Lastly, we observe that some (but not all) DL co-folding methods are highly dependent on the availability of diverse input MSAs to achieve high structural prediction accuracy, underscoring the need in future work to elucidate the impact of the availability of MSAs and protein language model embeddings on the training dynamics of biomolecular structure prediction methods. As a publicly available resource, PoseBench is flexible to accommodate new datasets and methods for protein-ligand docking and structure prediction.
@@ 4.2 Benchmark datasets
Name Type Source Size (Total # Ligands)
Astex Diverse Primary Ligand [16] 85
PoseBusters Benchmark Primary Ligand [12] 130/308
DockGen-E Primary Ligand 122
CASP15 Multi-Ligand 102 (across 19 complexes)
$\rightarrow$ 6 (13) single (multi)-ligand complexes
@@
PoseBusters Benchmark dataset. Version 2 of the the popular PoseBusters Benchmark dataset [12], which we adopt in this work, contains 308 recent primary ligand protein complexes deposited in the PDB from 2019 onwards. Accordingly, in contrast to Astex Diverse, this dataset can be considered a moderately difficult benchmark dataset for baseline methods, since many of its complexes do not directly overlap with the most commonly used PDB-based training data. Important to note is that, among all baseline methods, AF3 used the most recent PDB training data cutoff of September 30, 2021, which motivated us to report the results in Section 2.3 for only the subset of PoseBusters Benchmark complexes (n=130) deposited in the PDB after this date. Like Astex Diverse, for the PoseBusters Benchmark dataset, we used AF3 (and ESMFold) to predict the apo protein structures of each of its complexes and then performed our PyMOL-based structural binding site alignments.
DockGen-E dataset. The original DockGen dataset [13] contains 189 diverse primary ligand protein complexes, each representing a functionally distinct type of PLI binding pocket according to ECOD domain partitioning [19, 20]. Consequently, this dataset can be considered PoseBench’s most difficult primary ligand dataset to model since its PLI binding sites are distinctly uncommon compared to those frequently found in the training datasets of all baseline methods, though it is important to note that these original DockGen complexes were deposited in the PDB from 2019 onward, making this benchmarking dataset partially overlap with the training datasets of baseline DL co-folding methods such as AF3 and Chai-1. Nonetheless, in line with our initial hypotheses, the benchmarking results in Section 2 demonstrate that no baseline method can adequately predict the PLI binding sites and ligand poses represented by this bespoke subset of the PDB, suggesting that all baseline DL methods have yet to learn broadly applicable representations of protein-ligand binding.
@@ 4.4 Metrics
Lastly, we also measure the molecule validity rates of each predicted PLI complex pose using the PoseBusters software suite (i.e., PB-Valid) [12]. This suite runs several important chemical and structural sanity checks for each predicted pose including energy ratio inspection and geometric (e.g., flat ring) assertions which provide a secondary filter of accurate poses that are also chemically and structurally meaningful.
@@
[…]
where $\Pi(u,v)$ denotes the set of distributions on $\mathbb{R}\times\mathbb{R}$ whose marginals, $u$ and $v$, are on the first and second factors, respectively. To penalize a baseline method for producing non-native interaction types, we unify the bins in each histogram before converting them into 1D discrete representations. Namely, to perform this calculation, each PLI is first represented as a fingerprint tuple of $<$ligand type, amino acid type, interaction type$>$ as determined by the software tool ProLIF [26] and then grouped to count each type of tuple to form a histogram. As such, a lower PLIF-EMD value implies a better continuous agreement between predicted and native interaction histograms. PLIF-WM, derived from PLIF-EMD, assesses the Wasserstein matching (WM) score of a pair of PLIF histograms. Specifically, to obtain a more benchmarking-friendly score ranging from 0 to 1 (higher is better), we define PLIF-WM as
[…]
@@ 4.5 Baseline methods and experimental setup
As a representative algorithm for conventional protein-ligand docking, we pair AutoDock Vina (v1.2.5) [31] for molecular docking with P2Rank for protein-ligand binding site prediction [32] to form a strong conventional (blind) docking baseline (P2Rank-Vina) for comparison with DL methods. To represent DL docking methods, we include DiffDock-L [13] for docking with static protein structures and DynamicBind [9] for flexible docking. Lastly, to represent some of the latest DL co-folding methods, we include NeuralPLexer [10], RFAA [7], AF3 [11], and Chai-1 [17]. For interested readers, each method’s input and output data formats are described in Appendix F.
@@ 5 Related work
Protein-ligand structure prediction and docking. Shortly following the first wave of predictive methods for protein-ligand structure determination, DL methods such as DiffDock [8] demonstrated the utility of a new approach to this problem by reframing protein-ligand docking as a generative modeling task, whereby multiple ligand conformations can be generated for a particular protein target and rank-ordered using a predicted confidence score [34]. This approach has inspired many follow-up works offering alternative formulations of this generative approach to the problem [35, 36, 37, 38, 39, 9, 40, 41, 10, 42, 43, 13, 7, 44, 45, 46, 47, 11, 17, 48], with some of such follow-up works also being capable of accurately modeling protein flexibility upon ligand binding or predicting binding affinities to a high degree of accuracy.
Benchmarking efforts for protein-ligand complexes. In response to the large number of new methods that have been developed for protein-ligand structure prediction, recent works have introduced several new datasets and metrics with which to evaluate newly developed methods, with some of such benchmarking efforts focusing on modeling single-ligand protein interactions [49, 12, 50, 28, 51, 52, 53] and others specializing in the assessment of multi-ligand protein interactions [25]. One of the motivations for introducing PoseBench in this work is to bridge this gap by systematically assessing a selection of the latest (pocket-blind) structure prediction methods within both interaction regimes, using unbound (apo) protein structures with docking methods and challenging DL co-folding methods to predict full bioassemblies from primary sequences. Our benchmarking results in Section 2 demonstrate the relevance and utility of this comprehensive new evaluation suite for the future of protein-ligand modeling.
@@ 6 Data availability
The PoseBench datasets and benchmark results are available at https://zenodo.org/records/14629652 under a Creative Commons Attribution 4.0 International Public License, with further licensing discussed in Appendix A and detailed dataset documentation provided in Appendix D.
@@ Appendix D Documentation for datasets
Below, we provide detailed documentation for each dataset included in our benchmark, summarized in Table 1 of the main text. Each dataset is freely available for download from the benchmark’s accompanying Zenodo data record [54] under a CC-BY 4.0 license. In lieu of being able to create associated metadata for each of our macromolecular datasets using an ML-focused library such as Croissant [55] (due to file type compatibility issues), instead, we report structured metadata for our preprocessed datasets using Zenodo’s web user interface [54]. Note that, for all datasets, we authors bear all responsibility in case of any violation of rights regarding the usage of such datasets.
@@ D.1 Astex Diverse Set - Primary Ligand Docking
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.14629652) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/14629652/files/astex_diverse_set.tar.gz.
@@ D.2 PoseBusters Benchmark Set - Primary Ligand Docking
  • Composition The dataset consists of 308 primary ligand protein complexes deposited in the PDB in 2019 and after. As such, this dataset poses a moderate challenge for DL methods, since several of such methods were trained on data deposited before this cutoff date (notably except for Chai-1 and AF3 which used training cutoff dates of January 12, 2021 and September 30, 2021, respectively). For each of the dataset’s complexes, we obtained high-accuracy predicted protein structures using AF3. The accuracy of the AF3-predicted structures is measured in terms of their RMSD and TM-score compared to the corresponding crystal protein structures and is visualized in Figure 12. Notably, after alignment with the crystalized (holo) PLI binding pocket residues, 59.09% (53.25% with ESMFold) of the predicted structures have a global RMSD below 4 Å and TM-score above 0.7, indicating that most of the dataset’s proteins have a reasonably accurate predicted structure.
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.14629652) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/14629652/files/posebusters_benchmark_set.tar.gz.
@@ D.3 DockGen-E Set - Primary Ligand Docking
  • Collection To curate this dataset, we collected the original dataset’s protein and ligand binding pocket annotations for each complex introduced by Corso et al. [13]. Subsequently, we retrieved the corresponding first biological assembly listed in the PDB to obtain each PDB entry’s biologically relevant complex, filtering out complexes for which the first assembly could not be mapped to its original protein and ligand binding pocket annotation. This procedure left 122 biologically relevant assemblies remaining for benchmarking. Important to note is that these original DockGen complexes were deposited in the PDB from 2019 onward, making this benchmarking dataset partially overlap with the training datasets of multiple DL co-folding baseline methods such as NeuralPLexer, AF3, and Chai-1. Nonetheless, our benchmarking results in the main text demonstrate that baseline DL methods are challenged to find the correct (novel) binding pocket conformations represented by this dataset, suggesting that all baseline DL models have yet to learn truly comprehensive representations of protein-ligand binding.
@@
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.14629652) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/14629652/files/dockgen_set.tar.gz.
@@ D.4 CASP15 Set - Multi-Ligand Docking
  • Hosting Our preprocessed version of (the publicly available version of) this dataset (https://doi.org/10.5281/zenodo.14629652) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/14629652/files/casp15_set.tar.gz.
@@ E.1 Dataset protein-ligand interaction distributions
Overall, we find that the PDBBind 2020, Astex Diverse, and PoseBusters Benchmark datasets contain similar types and frequencies of interactions, with the PoseBusters Benchmark dataset containing slightly more hydrogen bond acceptors ($\sim$3 vs 1) and fewer van der Waals contacts ($\sim$5 vs 8) on average compared to the PDBBind 2020 dataset. However, we observe a more notable difference in interaction types and frequencies between the DockGen and CASP15 datasets and the three other datasets. Specifically, we find these two benchmark datasets contain a notably different quantity of hydrogen bond acceptors and donors (n.b., $\sim$40 for CASP15), van der Waals contacts ($\sim$200 for CASP15), and hydrophobic interactions ($\sim$2 for DockGen) on average. These dataset-level interaction disparities may partially explain the baseline-challenging DockGen benchmarking results reported in Section 2 of the main text.
@@ E.2 Baseline method protein-ligand interaction distributions
Intrigued by the dataset interaction patterns presented in Figure 15, here we further investigate the predicted PLIs produced by each baseline method for each evaluation dataset to study which DL methods can most faithfully reproduce the native distribution of PLIs within each dataset. Our results in Figures 16, 17, 18, and 19 suggest that AF3 demonstrates the best overall ability to recapitulate the crystalized PLIs observed within these datasets, in line with the PLIF-WM benchmarking results presented in Section 2 of the main text. Nonetheless, its predicted interaction distributions, in particular for the DockGen and CASP15 datasets, have much room for improvement, especially for more structured interactions such as hydrogen bonds.

### v6

@@ 3 Conclusions
In this work, we have introduced PoseBench, a unified, broadly applicable benchmark and toolkit for studying the performance of methods for protein-ligand docking and structure prediction. Benchmarking results with PoseBench, summarized in Table 1, suggest that DL co-folding methods generally outperform conventional and DL docking baselines yet remain challenged to predict complexes containing novel protein-ligand binding poses, with AF3 performing best overall when deep MSAs are available for a target protein (Chai-1 otherwise), regardless of the availability of homologous proteins. Further, we find that several DL methods face difficulties balancing the structural accuracy of their predicted poses with the chemical specificity of their induced protein-ligand interactions, highlighting that future methods may benefit from the introduction of physico-chemical loss functions or sampling techniques to bridge this performance gap. Lastly, we observe that some (but not all) DL co-folding methods are highly dependent on the availability of diverse input MSAs to achieve high structural prediction accuracy (e.g., AF3 but not Chai-1 or Boltz-1), underscoring the need in future work to elucidate the impact of the availability of MSAs and protein language model embeddings on the training dynamics of biomolecular structure prediction methods. As a publicly available resource, PoseBench is flexible to accommodate new datasets, methods, and analyses for protein-ligand docking and structure prediction.
@@ 4.2 Benchmark datasets
Astex Diverse DockGen-E PoseBusters Benchmark CASP15
Ligand Type Primary Primary Primary Multi
Source [16] [12]
Size (Total # Ligands) 85 122 130/308 102 (across 19 complexes) $\rightarrow$ 6 (13) single (multi)-ligand complexes
Training Data Homology High Moderate Low Low
Top-Ranked Method (w/ MSAs) AF3 AF3 AF3 AF3
Top-Ranked Method (w/o MSAs) Chai-1 Chai-1 Chai-1 NeuralPLexer
@@
PoseBusters Benchmark dataset. Version 2 of the the popular PoseBusters Benchmark dataset [12], which we adopt in this work, contains 308 recent primary ligand protein complexes deposited in the PDB from 2019 onwards. Accordingly, in contrast to Astex Diverse, this dataset can be considered a moderately difficult benchmark dataset for baseline methods, since many of its complexes do not directly overlap with the most commonly used PDB-based training data. Important to note is that, among all baseline methods, AF3 and Boltz-1 used the most recent PDB training data cutoff of September 30, 2021, which motivated us to report the results in Section 2.3 for only the subset of PoseBusters Benchmark complexes (n=130) deposited in the PDB after this date. Like Astex Diverse, for the PoseBusters Benchmark dataset, we used AF3 (and ESMFold) to predict the apo protein structures of each of its complexes and then performed our PyMOL-based structural binding site alignments.
DockGen-E dataset. The original DockGen dataset [13] contains 189 diverse primary ligand protein complexes, each representing a functionally distinct type of PLI binding pocket according to ECOD domain partitioning [20, 21]. Consequently, this dataset can be considered PoseBench’s most difficult primary ligand dataset to model since its PLI binding sites are distinctly uncommon compared to those frequently found in the training datasets of all baseline methods, though it is important to note that these original DockGen complexes were deposited in the PDB from 2019 onward, making this benchmarking dataset partially overlap with the training datasets of baseline DL co-folding methods such as Chai-1, Boltz-1, and AF3. Nonetheless, in line with our initial hypotheses, the benchmarking results in Section 2 demonstrate that no baseline method can adequately predict the PLI binding sites and ligand poses represented by this bespoke subset of the PDB, suggesting that all baseline DL methods have yet to learn broadly applicable representations of protein-ligand binding.
@@ 4.4 Metrics
Lastly, we also measure the molecule validity rates of each predicted PLI complex pose using the PoseBusters software suite (i.e., PB-Valid) [12]. This suite runs several important chemical and structural sanity checks for each predicted pose including energy ratio inspection and geometric (e.g., flat aliphatic ring) assertions which provide a secondary filter of accurate poses that are also chemically and structurally meaningful.
@@
$\text{PLIF-EMD}\ :=\ l_{1}(u,v)=\inf_{\pi\in\Pi(u,v)}\int_{\mathbb{R}\times\mathbb{R}}|x-y|d\pi(x,y),$ (3)
where $\Pi(u,v)$ denotes the set of distributions on $\mathbb{R}\times\mathbb{R}$ whose marginals, $u$ and $v$, are on the first and second factors, respectively. To penalize a baseline method for producing non-native interaction types, we unify the bins in each histogram before converting them into 1D discrete representations. Namely, to perform this calculation, each PLI is first represented as a fingerprint tuple of <<ligand type, amino acid type, interaction type>> as determined by the software tool ProLIF [29] and then grouped to count each type of tuple to form a histogram. As such, a lower PLIF-EMD value implies a better continuous agreement between predicted and native interaction histograms. PLIF-WM, derived from PLIF-EMD, assesses the Wasserstein matching (WM) score of a pair of PLIF histograms. Specifically, to obtain a more benchmarking-friendly score ranging from 0 to 1 (higher is better), we define PLIF-WM as
$\text{PLIF-WM}\ :=\ 1-\frac{\text{PLIF-EMD}-\text{PLIF-EMD}_{\text{min}}}{\text{PLIF-EMD}_{\text{max}}-\text{PLIF-EMD}_{\text{min}}},$ (4)
where $\text{PLIF-EMD}_{\text{min}}$ and $\text{PLIF-EMD}_{\text{max}}$ denote the minimum (best) and maximum (worst) values of PLIF-EMD, respectively. As a metric normalized relative to each collection of the latest baseline methods, PLIF-WM allows one to quickly identify which of the latest methods has the greatest capacity to produce realistic distributions of PLIs. As a practical note, we use SciPy 1.15.1 [33] to provide users of PoseBench with an optimized implementation of PLIF-EMD and thereby PLIF-WM.
@@ 4.5 Baseline methods and experimental setup
As a representative algorithm for conventional protein-ligand docking, we pair AutoDock Vina (v1.2.5) [34] for molecular docking with P2Rank for protein-ligand binding site prediction [35] to form a strong conventional (blind) docking baseline (P2Rank-Vina) for comparison with DL methods. To represent DL docking methods, we include DiffDock-L [13] for docking with static protein structures and DynamicBind [9] for flexible docking. Lastly, to represent some of the latest DL co-folding methods, we include NeuralPLexer [10], RFAA [7], Chai-1 [17], Boltz-1 [18] (versus Boltz-2 [36] for sake of time-split benchmarking validity), and AF3 [11]. For interested readers, each method’s input and output data formats are described in Appendix F.
@@ 5 Related work
Protein-ligand structure prediction and docking. Shortly following the first wave of predictive methods for protein-ligand structure determination, DL methods such as DiffDock [8] demonstrated the utility of a new approach to this problem by reframing protein-ligand docking as a generative modeling task, whereby multiple ligand conformations can be generated for a particular protein target and rank-ordered using a predicted confidence score [38]. This approach has inspired many follow-up works offering alternative formulations of this generative approach to the problem [39, 40, 41, 42, 43, 9, 44, 45, 10, 46, 47, 13, 7, 48, 49, 50, 51, 11, 17, 18, 52, 36], with some of such follow-up works also being capable of accurately modeling protein flexibility upon ligand binding or predicting binding affinities to a high degree of accuracy.
Benchmarking efforts for protein-ligand complexes. In response to the large number of new methods that have been developed for protein-ligand structure prediction, recent works have introduced several new datasets and metrics with which to evaluate newly developed methods [24], with some of such benchmarking efforts focusing on modeling single-ligand protein interactions [53, 12, 22, 31, 54, 55, 56] and others specializing in the assessment of multi-ligand protein interactions [28]. One of the motivations for introducing PoseBench in this work is to bridge this gap by systematically assessing a selection of the latest (pocket-blind) structure prediction methods within both interaction regimes, using unbound (apo) protein structures with docking methods and challenging DL co-folding methods to predict full bioassemblies from primary sequences. Our benchmarking results in Section 2 demonstrate the relevance and utility of this comprehensive new evaluation suite for the future of protein-ligand modeling.
@@ 6 Data availability
The PoseBench datasets and benchmark results are available at https://zenodo.org/records/16791095 under a Creative Commons Attribution 4.0 International Public License, with further licensing discussed in Appendix A and detailed dataset documentation (e.g., of AF3’s predicted protein structure accuracy) provided in Appendix D.
@@ Appendix D Documentation for datasets
Below, we provide detailed documentation for each dataset included in our benchmark, summarized in Table 1. Each dataset is freely available for download from the benchmark’s accompanying Zenodo data record [57] under a CC-BY 4.0 license. In lieu of being able to create associated metadata for each of our macromolecular datasets using an ML-focused library such as Croissant [58] (due to file type compatibility issues), instead, we report structured metadata for our preprocessed datasets using Zenodo’s web user interface [57]. Note that, for all datasets, we authors bear all responsibility in case of any violation of rights regarding the usage of such datasets.
@@ D.1 Astex Diverse Set - Primary Ligand Docking
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.16791095) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/16791095/files/astex_diverse_set.tar.gz.
@@ D.2 PoseBusters Benchmark Set - Primary Ligand Docking
  • Composition The dataset consists of 308 primary ligand protein complexes deposited in the PDB in 2019 and after. As such, this dataset poses a moderate challenge for DL methods, since several of such methods were trained on data deposited before this cutoff date (notably except for Chai-1 and AF3/Boltz-1 which used training cutoff dates of January 12, 2021 and September 30, 2021, respectively). For each of the dataset’s complexes, we obtained high-accuracy predicted protein structures using AF3. The accuracy of the AF3-predicted structures is measured in terms of their RMSD and TM-score compared to the corresponding crystal protein structures and is visualized in Figure 12. Notably, after alignment with the crystalized (holo) PLI binding pocket residues, 59.09% (53.25% with ESMFold) of the predicted structures have a global RMSD below 4 Å and TM-score above 0.7, indicating that most of the dataset’s proteins have a reasonably accurate predicted structure.
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.16791095) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/16791095/files/posebusters_benchmark_set.tar.gz.
@@ D.3 DockGen-E Set - Primary Ligand Docking
  • Collection To curate this dataset, we collected the original dataset’s protein and ligand binding pocket annotations for each complex introduced by Corso et al. [13]. Subsequently, we retrieved the corresponding first biological assembly listed in the PDB to obtain each PDB entry’s biologically relevant complex, filtering out complexes for which the first assembly could not be mapped to its original protein and ligand binding pocket annotation. This procedure left 122 biologically relevant assemblies remaining for benchmarking. Important to note is that these original DockGen complexes were deposited in the PDB from 2019 onward, making this benchmarking dataset partially overlap with the training datasets of multiple DL co-folding baseline methods such as NeuralPLexer, Chai-1, Boltz-1, and AF3. Nonetheless, our benchmarking results in Section 2.2 demonstrate that baseline DL methods are challenged to find the correct (novel) binding pocket conformations represented by this dataset, suggesting that all baseline DL models have yet to learn truly comprehensive representations of protein-ligand binding.
@@
  • Hosting Our preprocessed version of the dataset (https://doi.org/10.5281/zenodo.16791095) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/16791095/files/dockgen_set.tar.gz.
@@ D.4 CASP15 Set - Multi-Ligand Docking
  • Hosting Our preprocessed version of (the publicly available version of) this dataset (https://doi.org/10.5281/zenodo.16791095) can be downloaded from the benchmark’s Zenodo data record at https://zenodo.org/records/16791095/files/casp15_set.tar.gz.
@@ E.1 Dataset protein-ligand interaction distributions
Overall, we find that the PDBBind 2020, Astex Diverse, and PoseBusters Benchmark datasets contain similar types and frequencies of interactions, with the PoseBusters Benchmark dataset containing slightly more hydrogen bond acceptors ($\sim$3 vs 1) and fewer van der Waals contacts ($\sim$5 vs 8) on average compared to the PDBBind 2020 dataset. However, we observe a more notable difference in interaction types and frequencies between the DockGen and CASP15 datasets and the three other datasets. Specifically, we find these two benchmark datasets contain a notably different quantity of hydrogen bond acceptors and donors (n.b., $\sim$40 for CASP15), van der Waals contacts ($\sim$200 for CASP15), and hydrophobic interactions ($\sim$2 for DockGen) on average. These dataset-level interaction disparities may partially explain the baseline-challenging DockGen benchmarking results reported in Section 2.
@@ E.2 Baseline method protein-ligand interaction distributions
Intrigued by the dataset interaction patterns presented in Figure 15, here we further investigate the predicted PLIs produced by each baseline method for each evaluation dataset to study which DL methods can most faithfully reproduce the native distribution of PLIs within each dataset. Our results in Figures 16, 17, 18, and 19 suggest that AF3 demonstrates the best overall ability to recapitulate the crystalized PLIs observed within these datasets, in line with the PLIF-WM benchmarking results presented in Section 2. Nonetheless, its predicted interaction distributions, in particular for the DockGen and CASP15 datasets, have much room for improvement, especially for more structured interactions such as hydrogen bonds.
@@ F.1 Input and output formats
    5. e)
      Boltz-1 is an open-source co-folding model (akin to AF3 and Chai-1) trained on exhaustive PDB crystal structures and AlphaFold 2-predicted distillation structures along with AF3-based training protocols. By default, the model receives as its inputs a (multi-chain) protein sequence and (fragment) ligand SMILES strings, with paired MSAs yet no template structures provided (as is its default setting). The method then produces multiple rank-ordered protein-ligand bound structure conformations for each input complex, using the method’s intrinsic AF3-like ranking quantity for rank-ordering. Note that, for the sake of benchmarking consistency (n.b., versus Chai-1), Boltz-1 uses standardized (taxonomy-paired) multiple sequence alignments akin to those used by AF3 in all benchmarking experiments. Note that, by default, we evaluate Boltz-1 with its inference-time potential functions enabled (i.e., Boltz-1x), making its predictions slightly slower but considerably more physically plausible overall (according to the PoseBusters software suite [60]), and we do not evaluate Boltz-2 (trained on PDB data deposited up to June 1, 2023) to maintain the validity of our benchmark’s time-splits.